### v1

# MLLM4PUE: Toward Universal Embeddings in Computational Pathology through Multimodal LLMs

Abstract
Pathology plays a critical role in diagnosing a wide range of diseases, yet existing approaches often rely heavily on task-specific models trained on extensive, well-labeled datasets. These methods face sustainability challenges due to the diversity of pathologies and the labor-intensive nature of data collection. To address these limitations, we highlight the need for universal multimodal embeddings that can support multiple downstream tasks. Previous approaches often involve fine-tuning CLIP-based models, which handle images and text separately, limiting their ability to capture complex multimodal relationships. Additionally, these models are evaluated across diverse datasets without a unified benchmark for assessing multimodal embeddings in pathology. To address these challenges, we propose MLLM4PUE, a novel framework that leverages Multimodal Large Language Models (MLLMs) to generate Pathology Universal Embeddings. The MLLM4PUE framework not only facilitates robust integration of images and text but also enhances understanding and fusion capabilities across various tasks. We further introduce the Pathology Multimodal Embedding Benchmark (PMEB), a comprehensive benchmark designed to assess the quality of pathology multimodal embeddings. PMEB comprises 15 original tasks drawn from 14 datasets, organized into three meta-tasks: retrieval, classification, and composed retrieval. Experimental results demonstrate the superiority of MLLM4PUE, illustrating MLLM-based models can effectively support a wide range of downstream tasks and unify the research direction for foundation models in pathology.
1 Introduction
Figure 1: Performance comparison across benchmark datasets, showing that our proposed MLLM4PUE achieves state-of-the-art (SOTA) results across all tasks. Weight F1 (wF1) is reported for classification task. For the retrieval and composed retrieval task, we report recall@5 scores. This highlights the advanced multimodal fusion capabilities of MLLM4PUE.
Pathology remains the gold standard for diagnosing a wide range of diseases [42, 32, 8]. With rapid advancements in artificial intelligence (AI), computational pathology has emerged as a promising field, using AI to address challenges across various tasks such as cancer diagnosis [29, 39], metastasis detection [44, 22], mutation prediction [37, 9, 33], and survival analysis [41, 40, 45]. These achievements, however, rely on task-specific models trained on large, well-labeled datasets. Given the hundreds of tumor types cataloged in the WHO classification111tumourclassification.iarc.who.int/ and the labor-intensive nature of data collection and annotation [12], developing separate, high-performing models for each pathology task is impractical, especially for rare diseases with limited data [24, 30]. Furthermore, while many efforts focus on image data, pathology often involves natural language, including diagnostic reports and research publications [6, 7, 31, 14]. Thus, there is a growing need to shift toward task-agnostic models that generate universal multimodal embeddings capable of supporting a broad range of pathology tasks.
In natural image processing, recent studies such as CLIP [34] have demonstrated remarkable success by aligning visual and language embeddings through contrastive learning, achieving strong performance across diverse downstream tasks. Building on these advances, several studies [30, 16, 38, 17, 31] have adapted CLIP-based models for pathology, training on extensive image-text pairs from public sources to produce multimodal embeddings that enable zero-shot learning across various pathology tasks. However, these methods encounter two significant challenges. First, CLIP [34] processes images and text independently, limiting its ability to capture complex interactions between modalities [19]. Second, evaluations of these models have mostly centered on isolated tasks, such as classification or retrieval [30, 38, 17], which separate image and text components and do not fully assess multimodal integration capabilities. Moreover, inconsistencies in datasets and sample selection criteria across studies lead to non-reproducible and incomparable results, emphasizing the need for a more integrated approach and a standardized benchmarking framework to advance multimodal embeddings in pathology.
To tackle the first challenge, we propose MLLM4PUE, a framework that uses a Multimodal Large Language Model (MLLM) backbone to generate universal multimodal embeddings for pathology. To the best of our knowledge, this is the first use of an MLLM backbone in this field, representing a significant advancement over previous CLIP-based methods. Unlike CLIP [34], which processes images and text separately, MLLM4PUE employs a transformer-based architecture to fully integrate these modalities, allowing it to learn from complex image-text relationships. This fusion enhances the model’s adaptability and effectiveness across a range of pathology tasks, making it particularly suited for generalizing to diverse pathology scenarios [28, 27].
To address the second challenge, we introduce the Pathology Multimodal Embedding Benchmark (PMEB) to establish a standardized evaluation platform for multimodal embeddings in pathology. This benchmark encompasses three meta-task categories: retrieval, composed retrieval, and classification. To ensure the tasks effectively assess multimodal embeddings, we have reformulated them to better suit this purpose. Each task provides the model with prompts to process a query and identify the correct target from a set of candidates. These queries and targets can consist of images, text, or a combination of both, enabling a flexible and comprehensive evaluation of multimodal capabilities. PMEB offers a standardized and reproducible approach for assessing multimodal embeddings in pathology. Overall, our contributions are summarized as follows:
  • We introduce MLLM4PUE, the first framework utilizing MLLMs to generate universal multimodal embeddings in pathology. This framework integrates multimodal information within a unified model, converting any combination of images and text into robust embeddings.
  • We provide PMEB, a comprehensive benchmark for evaluating multimodal embeddings in pathology, encompassing a broad range of tasks to support robust and generalizable assessment of these embeddings.
  • MLLM4PUE effectively represents multimodal information, consistently surpassing the performance of all baseline models across all evaluated tasks, as demonstrated by the quantitative results in Fig. 1.
  • Our approach offers new insights into computational pathology by adopting MLLMs for universal embeddings. Future research can build on this work to unify and streamline the development of models for diverse pathology tasks.
Figure 2: Overview of the MLLM4PUE framework. (a) Illustration of pathology image-text contrastive learning with our proposed method. With differents prompts, MLLM can generate embeddings for multimodal inputs. Image and text embeddings are aligned using a contrastive loss. (b) Zero transfer for classification. (c) Zero transfer for composed retrieval.
2 Related work
2.1 Multimodal embedding learning
Creating effective multimodal embeddings remains a challenging area of research. The CLIP model [34] addresses this by employing separate encoders for images and text, aligning them in a common space through contrastive learning. This approach has influenced many subsequent models, such as BLIP [26] and CoCa [43]. While these models excel in a variety of tasks, the use of distinct encoders restricts their capacity to fully integrate visual and textual data, which is a key requirement for tasks that involve combined visual-language inputs, such as search queries involving both images and text.
With the rise of large language models (LLMs), Multimodal Large Language Models (MLLMs) have extended LLMs to process multiple data modalities, achieving notable progress in understanding and reasoning across diverse input types [28, 27, 25]. Although MLLMs exhibit strong capabilities in interpreting multimodal content and following complex instructions, there is still limited research on their application in creating effective multimodal embeddings. The recent E5-V [19] model fine-tunes MLLMs with text-based data to improve embedding quality. Building on this, we have adapted the approach for pathology by using domain-specific image-text data, resulting in substantial performance improvements.
2.2 Multimodal embeddings for computational pathology
In computational pathology, recent studies have adopted the CLIP [34] architecture and training methodology, leveraging paired image-text data within a contrastive learning framework to align similar image-text embeddings while separating dissimilar ones [30, 16, 38, 17, 18]. For instance, PLIP [16] fine-tunes CLIP on large collections of image-text pairs sourced from platforms like Twitter and other public datasets. Similarly, PathCLIP [38] and QuiltNet [17] scale up pathology-specific data for fine-tuning, while CONCH [31] employs over 1.17 million image-caption pairs, making it the largest dataset used for training to date.
Despite these advancements, these methods fine-tune either CLIP [34] or CoCa [43] using pathology image-caption pairs and rely on separate encoders for processing images and text. In contrast, we propose an MLLM-based approach to capture universal multimodal embeddings for pathology by integrating image and text data within a unified model. This integrated framework enhances performance across a wide range of computational pathology tasks, offering a more holistic and effective solution.
3 Methodology
In this section, we first introduce the Pathology Multimodal Embedding Benchmark (PMEB), a comprehensive benchmark for evaluating multimodal embeddings in diverse pathology tasks. Then, we present the MLLM4PUE framework, which integrates multimodal embeddings, contrastive learning, and zero-shot transfer methods to align and adapt pathology image-text data effectively, as illustrated in Fig. 2.
3.1 Pathology multimodal embedding enchmark
Meta-Task Data source Query Target Original Tasks Original samples Selected samples
Retrieval Arch-book [13] I T Image-text retrieval 1,306 1,306
Arch-book [13] T I Text-image retrieval 1,306 1,306
Arch-pubmed [13] I T Image-text retrieval 1,923 1,923
Arch-pubmed [13] T I Text-image retrieval 1,923 1,923
Composed Quilt-VQA [35] I+T T Histopathology VQA 985 724
retrieval Quilt-VQA-RED [35] I+T T Histopathology VQA 335 252
Classification Bach [2] I T Breast tissue 399 399
CRC-100k [20] I T Colorectal cancer 100,000 1,500
Digestpath [10] I T Signet ring cell 455 455
Digestpath [10] I T Colonoscopy tissue 660 660
Kathercolon [21] I T Colorectal cancer tissue 7,180 1,500
LC25000 [4] I T Colon adenocarcinoma 10,000 1,500
LC25000 [4] I T Lung adenocarcinoma 15,000 1,500
Osteo [3] I T Osteosarcoma 1,144 1,144
Renalcell [5] I T Renal tissue 36,687 1,500
Sicap [36] I T Gleason pattern 12,081 1,500
Skincancer [23] I T Skin cancer tissue 129,369 1,500
Wsss4luad [15] I T LUAD tissue 4,693 1,500
Table 1: The statistics of our PMEB benchmark: 15 original tasks from 14 datasets. I and T denote image and text, respectively.
We propose PMEB (Pathology Multimodal Embedding Benchmark), a comprehensive benchmark designed to evaluate multimodal embeddings in pathology across a diverse range of tasks. These tasks include cancer tissue classification [2, 4, 23], Gleason pattern grading [36], and cell identification [10], among other tasks. PMEB comprises 15 original tasks from 14 datasets, organized into three meta-tasks: retrieval, classification, and composed retrieval. All tasks are reformulated to assess multimodal embeddings, where the model receives an instructional prompt and a query, which may consist of text, images, or both, and selects the correct target from a set of options.
In the retrieval task, the query can be either an image or text, with the target being the corresponding paired text or image. For the classification task, queries are images with targets as text labels representing various classes. The composed retrieval task involves queries that combine an image with text formatted as a question, with the target being the expected answer. This composed retrieval task, specifically designed for multimodal embedding evaluation, incorporates integrated visual and language inputs and has not been applied in previous methods. Detailed dataset statistics are provided in Table  1.
Due to the large size of certain classification datasets, we have standardized their sizes. For datasets containing more than 1,500 samples, we employ the CONCH model [31] to calculate similarity scores between each image and its corresponding class label text. We then select the top 500 images with the highest scores, the bottom 500 with the lowest scores, and a random sample of 500 from the remaining data, ensuring the original category distribution is preserved. For the composed retrieval task, we select only open-ended questions from the data source, where answers are more complex than a simple “yes” or “no”. Additionally, since there is an overlap between questions and answers in the VQA data source, we utilize ChatGPT [1] for data cleaning to ensure clarity and accuracy in task evaluation. Further details and examples from each dataset within PMEB are provided in Supplementary Materials.
Figure 3: Illustration of the MLLM4PUE framework for generating multimodal embeddings using an MLLM. The image is processed through an Image Encoder and projection layer, with a prompt guiding the MLLM to distill the image content into a single-word embedding. The final embedding $h$ is extracted from the last token as the unified multimodal representation.
3.2 Multimodal Embeddings with MLLM
We first introduce how our proposed MLLM4PUE leverages MLLMs for universal multimodal embedding learning. Unlike CLIP, which directly generates embeddings for multimodal inputs, MLLMs are inherently generative models and require a different approach to produce embeddings. Inspired by E5-V [19], our solution uses a prompt-based strategy specifically designed to guide MLLMs in embedding multimodal data.
[…]
The prompt design incorporates two components: first, the word Summarize directs the MLLM to distill the information of the multimodal input. The second part, in one word:, instructs the MLLM to compress this distilled information into a single token, thereby facilitating a unified multimodal embedding. Once the extended input $q_{input}$ is constructed, it is processed by the MLLM $f_{\varphi}$ to generate embeddings: $h=f_{\varphi}(q_{input})$, where $h$ is the embedding of last token obtained from the MLLM output, as illustrated in Fig. 3. This approach tailors the generation capabilities of MLLMs to effectively produce and utilize multimodal embeddings, expanding their application potential beyond traditional generative outputs.
3.3 Pathology Contrastive Learning
Next, we describe the approach for projecting pathological data into a latent embedding space using conservative learning. As illustrated in Fig. 2, given a batch of $N$ paired image and text samples $\{(v_{n},t_{n})\}_{n=1,\dots,N}$ and MLLM $f_{\varphi}$, for each sample $(v_{i},t_{i})$, we feed them into $f_{\varphi}$ to get the image and text embeddings $(h_{vi},h_{ti})$ as described in Sec. 3.2.
The objective is to build a latent space for pathology image-text embeddings that maximizes the similarity between embeddings of the paired samples in the batch, while minimizing the similarity between embeddings of the $N-1$ incorrect pairs. We optimize an infoNCE contrastive Loss $\mathcal{L}$ to train our model.
$\mathcal{L}=-(\log\frac{e^{cos(h_{vi},h_{ti})/\tau}}{\sum_{j=1}^{n}e^{cos(h_{vi},h_{tj})/\tau}}+\log\frac{e^{cos(h_{ti},h_{vi})/\tau}}{\sum_{j=1}^{n}e^{cos(h_{ti},h_{vj})/\tau}}),$ (1)
where $\tau$ is a temperature parameter and $cos(h_{vi},h_{ti})$ and $cos(h_{ti},h_{vi})$ represent the cosine similarity in both directions in contrastive learning.
3.4 Zero-shot Transfer Evaluation
Our model, trained primarily on retrieval tasks focused on image-text pairs, requires adaptation to handle zero-shot classification and composed retrieval tasks. The details of how we modify the model for these different zero-shot downstream tasks are introduced as follows.
For zero-shot classification, we adopt a prompt-based method inspired by the CLIP model [34]. In this approach, each class name is expanded into a sentence using a specific template. For instance, the class name “Colon adenocarcinoma” is converted into the sentence “An H&E image of Colon adenocarcinoma” using the template “An H&E image of { }.” We apply this method to create sentences for all class names. Our model then computes embeddings for these sentences and the test images, extracting the last token of the MLLM output as described in Section 3.2. The similarity between these embeddings is calculated as outlined in Section 3.3, and test image labels are assigned based on the highest similarity scores. To account for variability in templates, we randomly select one from the 21 templates in CONCH [31] and repeat this process 100 times for each dataset. The templates are exhibited in Supplementary Materials.
[…]
Task Model Arch-PubMed [13] Arch-book [13]
Recall@5 Recall@10 Recall@50 Recall@5 Recall@10 Recall@50
i2t E5-V [19] 0.006 0.011 0.052 0.009 0.018 0.067
PLIP [16] 0.037 0.067 0.185 0.096 0.152 0.393
MLLM4PUE-O 0.105 0.160 0.366 0.185 0.264 0.575
PathCLIP [17] 0.275 0.388 0.680 0.152 0.234 0.482
MLLM4PUE-P 0.372 0.495 0.782 0.192 0.283 0.603
QuiltNet [17] 0.069 0.111 0.273 0.116 0.168 0.384
MLLM4PUE-Q 0.122 0.177 0.407 0.182 0.248 0.502
t2i E5-V [19] 0.006 0.011 0.052 0.009 0.018 0.067
PLIP [16] 0.037 0.067 0.181 0.112 0.164 0.419
MLLM4PUE-O 0.114 0.173 0.385 0.193 0.280 0.600
PathCLIP [17] 0.236 0.348 0.630 0.137 0.196 0.445
MLLM4PUE-P 0.297 0.399 0.688 0.185 0.277 0.555
QuiltNet [17] 0.056 0.092 0.237 0.100 0.152 0.389
MLLM4PUE-Q 0.135 0.193 0.433 0.210 0.302 0.584
Table 2: Performance comparison with baseline methods on two retrieval datasets, reporting recall metrics at different thresholds. Task i2t and t2i denote image-to-text and text-to-image retrieval, respectively. MLLM4PUE-O, MLLM4PUE-P, and MLLM4PUE-Q denote our model pre-trained on Openpath [16], PathCap [38], Quilt1M [17], respectively. Bold values indicate the best performance.
4 Experiments
Figure 4: Comparison of zero-shot patch classification performance across different models, measured by weighted F1 metric. Each box represents 100 data points, with each point indicating the performance on one template. The upper, middle, and lower lines of each box denote the first, median, and third quartiles, respectively. MLLM4PUE-O, MLLM4PUE-P, and MLLM4PUE-Q denote our model pre-trained on Openpath [16], PathCap [38], Quilt1M [17], respectively.
4.1 Training Datasets
We leverage three public datasets, Openpath [16], PathCap [38] and Quilt1M [17], to construct a robust training foundation for pathology image-text pairs.
The Openpath dataset [16] is assembled by gathering data from Twitter and LAION using popular pathology hashtags. Since PLIP[16] provided only Twitter IDs, we retrieve the corresponding data from Twitter and LAION using these IDs, resulting in a dataset of 138,874 image-text pairs after denoising. The PatchCap dataset [38] comprises 207K high-quality samples from authoritative sources, and we use all provided image-text samples. Similarly, the Quilt1M dataset [17] aggregates pathology image-text pairs from several public sources. As many images in Quilt1M are associated with multiple captions, we concatenate the captions for each image to create comprehensive entries. Through these careful preparations, we compile an extensive training dataset consisting of 593,838 image-text pairs.
4.2 Baselines
We evaluate the performance of our proposed MLLM4PUE, against several recent SOTA approaches across various downstream tasks. Specifically, we compare with three currently developed CLIP-based methods: PLIP [16], PathCLIP [38], and QuiltNet [17]. We also compare our method with E5-V [19], a contemporary model that leverages vision-language models for multimodal embedding tasks. It proposes a single-modality training approach, where the model is trained exclusively on text pairs.
4.3 Implementation Details and Metrics
We use LLaVA-NeXT [25] as our backbone MLLM. To save the GPU memory, we use QLoRA [11] and gradient checkpointing with DeepSpeed ZeRO-2, and all the images are resized to 336 × 336 pixels. The temperature is set to 0.02. All experiments are run on six H100 GPUs.
To evaluate the performance of retrieval and composed retrieval tasks, we use the Recall@K metric, which quantifies the proportion of relevant items contained within the top K retrieved results. For classification tasks, due to the imbalanced distribution of classes, we use the weighted F1 (wF1) score to assess performance. The weighted F1 is calculated by averaging the F1 scores for each class, with each class’s score weighted by its frequency of occurrence.
5 Results
Model Quilt-VQA [35] Quilt-VQA-RED [35]
Recall@5 Recall@10 Recall@50 Recall@5 Recall@10 Recall@50
E5-V [19] 0.214 0.323 0.495 0.351 0.441 0.668
PLIP [16] 0.191 0.224 0.349 0.250 0.286 0.500
MLLM4PUE-O 0.239 0.334 0.583 0.373 0.460 0.790
PathCLIP [17] 0.192 0.242 0.413 0.262 0.337 0.603
MLLM4PUE-P 0.218 0.329 0.541 0.363 0.468 0.790
QuiltNet [17] 0.180 0.209 0.336 0.222 0.274 0.460
MLLM4PUE-Q 0.475 0.598 0.823 0.675 0.754 0.948
Table 3: Zero-shot composed retrieval performance on the Quilt-VQA and Quilt-VQA-RED datasets, with Recall metrics reported at various thresholds. MLLM4PUE-O, MLLM4PUE-P, and MLLM4PUE-Q denote our model pre-trained on Openpath [16], PathCap [38], Quilt1M [17], respectively. Bold values indecate the best performance.
Prompt-tuning Prompt Arch-PubMed [13] Arch-book [13]
i2t t2i i2t t2i
R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10
no no prompt 0.086 0.135 0.087 0.137 0.147 0.213 0.149 0.226
no “in one word” 0.092 0.137 0.095 0.142 0.152 0.236 0.158 0.233
yes no prompt 0.091 0.139 0.091 0.141 0.152 0.233 0.161 0.246
no “in one word” 0.094 0.143 0.102 0.149 0.157 0.243 0.165 0.258
MLLM4PUE 0.098 0.155 0.109 0.162 0.169 0.253 0.177 0.274
Table 4: Impact of prompt-tuning on two retrieval datasets. All models are trained on the Openpath dataset [16] for 1000 steps. Task i2t and t2i denote image-to-text and text-to-image retrieval. R@5 and R@10 represent Recall@5 and Recall@10, respectively. Bold values indicate the best performance.
Model Quilt-VQA [13] Quilt-VQA-RED [13]
Recall@5 Recall@10 Recall@50 Recall@5 Recall@10 Recall@50
QuiltNet [17] only question 0.170 0.189 0.275 0.187 0.234 0.440
add embedding 0.180 0.209 0.336 0.222 0.274 0.460
MLLM4PUE only question 0.291 0.362 0.539 0.429 0.488 0.694
add embedding 0.394 0.500 0.746 0.520 0.619 0.909
prompt 0.475 0.598 0.823 0.675 0.754 0.948
Table 5: Ablation study of image and question embedding fusion methods on composed retrieval task. All models are trained on the Quilt1M [17] dataset. Recall metrics are reported. Bold values indicate the best performance.
5.1 Zero-shot Cross-modal Retrieval
We first evaluate our methods on the retrieval task, where the query is an image or text, with the target being the corresponding paired text or image. Recall@k metric is employed to assess the model’s ability to retrieve relevant pathology image-text pairs among top k results. Our MLLM4PUE models are pretrained on Openpath [16], PatchCap [38], and Quilt1M [17]. As shown in Table 2, our models consistently outperform the baseline models, especially at higher recall values. Notably, in the image-to-text (i2t) task on Arch-book dataset [13], our MLLM4PUE trained on Openpath dataset [16] achieves a Recall@5 of 0.185, indicating an approximate 9% improvement over PLIP [16]. Moreover, our PathCap-pretrained model surpasses the performance of PathCLIP [38] on both benchmark datasets. A similar enhancement is observed in the text-to-image (t2i) retrieval task. For instance, on the Arch-PubMed dataset [13], our model achieves a Recall@10 of 0.193, marking a 10% increase in performance. These results collectively demonstrate that our framework excels at capturing multimodal embeddings more effectively than CLIP-based models with the same training data. Furthermore, it is important to note that the E5-V model [19], although sharing the same underlying architecture as ours, yields significantly lower recall scores across all tasks and datasets due to the lack of fine-tuning with pathology-specific data.
5.2 Zero-shot Patch Classification
We then evaluate the performance of zero-shot classification on patch-level pathology images. Specifically, each query is an image, and the target is an extended sentence of the class name, constructed using a template outlined in sec 3.4. Following CONCH [31], we conduct 100 experimental trials for each dataset, randomly selecting one template from a set of 21 templates for each trial. The detailed descriptions are available in the Supplementary Material. We present the distribution of weighted F1 scores (wF1) from these 100 experiments in Fig. 4, where the top, middle, and bottom lines of each box depict the first quartile, median, and third quartile of the wF1 score distribution, respectively. Each point in the figure indicates a single template’s performance.
Fig. 4 illustrates that our MLLM4PUE consistently surpasses baseline models trained on identical datasets. For instance, in the “DigestpathSignet” task with the Openpath dataset [16], which focuses on identifying signet ring cells, MLLM4PUE-O achieves a median wF1 score of 0.665, marking an improvement of over 30% compared to PLIP’s median of 0.329. When the training data changes to PathCLIP [38], MLLM4PUE-P also shows considerable advancements, achieving a median score of 0.854 compared to PathCLIP’s 0.361. This implies a performance increase of roughly 50%. Similarly, using the Quiltnet dataset [17], MLLM4PUE-Q excels in the “Sicap” dataset, reaching a median wF1 of 0.402, far exceeding Quiltnet’s 0.164. In the “Osteo” dataset, MLLM4PUE-Q achieves a median wF1 score of 0.671, significantly outclassing Quiltnet with an improvement exceeding 40%. Moreover, when compared with E5V, MLLM4PUE demonstrates superior performance, delivering higher wF1 scores even with the same architectural foundation. These results confirm the superiority of MLLM4PUE in enhancing zero-shot classification performance through effective multimodal data integration, establishing it as a robust choice for various pathology image classification tasks.
5.3 Zero-shot Composed Retrieval
We further evaluate the performance of composed retrieval on two datasets, where the task involves using an image and a question as the query to retrieve the correct answer. The Recall@k metric is employed to assess how well the model can identify the correct answer within the top k candidates. As shown in Table 3, our proposed MLLM4PUE demonstrates significant improvements compared to baseline methods, even when trained with the same pre-training data. Specifically, on the Quilt-VQA dataset [35], MLLM4PUE achieves an increase in Recall@5 by approximately 4%, 3%, and 30% over PLIP [16], PathCLIP [38], and QuiltNet [17], respectively. Similarly, on the Quilt-RED-VQA dataset [35], MLLM4PUE pre-trained on Quilt1M [17], has achieved a Recall@5 of 0.675, which represents the best performance among all baseline methods. Interestingly, the E5-V model [19] exhibits superior performance compared to all CLIP-based models, despite lacking fine-tuning on pathology-specific data. This demonstrates the robust capacity of MLLM backbone in multimodal embedding fusion, highlighting its proficiency in handling diverse data inputs effectively.
5.4 Ablation studies
5.4.1 Impact of Prompts
To assess the impact of different prompt strategies, we conduct a comparative analysis between our MLLM4PUE and two alternative approaches: no prompt - where images or text are directly input into the model without prompts, and no “in one word” - employing a similar prompt to our method but omitting the phrase “in one word”. Table 4 presents the performance of these methods, both with and without fine-tuning. All models are trained on Openpath dataset [16] for 1000 steps. For models employing fine-tuning, each method is adjusted with its respective prompts. In contrast, the no fine-tuning scenario involves simply altering the prompt for the already finely-tuned MLLM4PUE model. Our method shows significant improvements on zero-shot retrieval tasks compared to the no prompt and no “in one word”. For the Arch-PubMed dataset, MLLM4PUE attains Recall@5 scores of 0.098 and 0.109 for image-to-text and text-to-image tasks, respectively, marking the best performance in those categories. Similarly, in the Arch-book dataset, MLLM4PUE achieves the top Recall@5 and Recall@10 scores, indicating superior retrieval effectiveness. These results highlight the significant enhancements provided by our prompt, emphasizing its role in facilitating more accurate zero-shot retrieval in complex datasets by better aligning multimodal inputs.
5.4.2 Fusion Method in Composed Retrieval
This experiment is designed to highlight the effectiveness of our MLLM4PUE in modality fusion while exploring the limitations of CLIP in handling composed retrieval tasks. As shown in Table 5, the results reveal that QuiltNet [17], a CLIP-based model, shows little difference in performance between using only the question text and integrating image embeddings. This indicates that CLIP’s method lacks robust modality integration, as the inclusion of visual information fails to significantly enhance its effectiveness. Conversely, our model, MLLM4PUE, demonstrates substantial improvements by effectively fusing image and text information. MLLM4PUE outperforms QuiltNet in every setting. The true strength of our approach becomes evident when employing a prompt-based method, leading to significantly higher recall scores. Specifically, it achieves Recall@5 values of 0.475 on Quilt-VQA and 0.675 on Quilt-VQA-RED, representing approximately an 8% and 12% improvement over the simple addition method, respectively, and an 18% and 25% improvement over using the question text alone.
These findings clearly indicate that our model’s ability to integrate both visual and textual modalities results in significant enhancements in retrieval accuracy. This surpasses both mere modality addition and the question-only configuration, highlighting the effectiveness of our modality fusion technique for composed retrieval tasks. Our approach provides advantages that cannot be achieved through separate handling or limited integration of text and image data.
6 Conclusion
In conclusion, our work represents a significant advancement in different downstream tasks through MLLM4PUE, a novel framework leveraging MLLMs to achieve universal multimodal embeddings. By effectively integrating image and text data, our model addresses the limitations of previous CLIP-based methods. We also introduce PMEB, which provides a standardized and comprehensive benchmark for evaluating multimodal capabilities, promoting reproducibility and comparability in the field. This study also paves the way for future research to utilize MLLM-based models to support a wide range of downstream tasks and unify the research direction for foundation models in pathology.

### v2

@@
-MLLM4PUE: Toward Universal Embeddings in Computational Pathology through Multimodal LLMs
+Toward Universal Embeddings in Digital Pathology through Multimodal LLMs
 Abstract
@@
 To address these limitations, we highlight the need for universal multimodal embeddings that can support multiple downstream tasks.
-Previous approaches often involve fine-tuning CLIP-based models, which handle images and text separately, limiting their ability to capture complex multimodal relationships.
-Additionally, these models are evaluated across diverse datasets without a unified benchmark for assessing multimodal embeddings in pathology.
-To address these challenges, we propose MLLM4PUE, a novel framework that leverages Multimodal Large Language Models (MLLMs) to generate Pathology Universal Embeddings.
-The MLLM4PUE framework not only facilitates robust integration of images and text but also enhances understanding and fusion capabilities across various tasks.
-We further introduce the Pathology Multimodal Embedding Benchmark (PMEB), a comprehensive benchmark designed to assess the quality of pathology multimodal embeddings.
-PMEB comprises 15 original tasks drawn from 14 datasets, organized into three meta-tasks: retrieval, classification, and composed retrieval.
-Experimental results demonstrate the superiority of MLLM4PUE, illustrating MLLM-based models can effectively support a wide range of downstream tasks and unify the research direction for foundation models in pathology.
+Previous approaches involve fine-tuning CLIP-based models, which handle images and texts separately, limiting their ability to capture complex multimodal relationships.
+Additionally, these models are evaluated across diverse datasets without a unified benchmark.
+In this paper, we explore the possibility of applying Multimodal Large Language Models (MLLMs) to generate pathology universal embeddings to address these challenges.
+Our contributions can be summarized in the following aspects: 1) We propose MLLM4PUE, a novel framework that leverages MLLMs to generate embeddings for various pathology downstream tasks.
+2) We further introduce the Pathology Multimodal Embedding Benchmark (PMEB), a comprehensive benchmark designed to assess the quality of pathology multimodal embeddings, which comprises 16 original tasks drawn from 15 datasets.
+3) Extensive experimental results demonstrate the superiority of MLLM4PUE, illustrating MLLM-based models can effectively support a wide range of downstream tasks and unify the research direction for foundation models in pathology.
 1 Introduction
-Figure 1: Performance comparison across benchmark datasets, showing that our proposed MLLM4PUE achieves state-of-the-art (SOTA) results across all tasks.
+Figure 1: Performance comparison across benchmark datasets, showing that our proposed MLLM4PUE achieves baseline methods across most of tasks.
 Weight F1 (wF1) is reported for classification task.
@@
 This highlights the advanced multimodal fusion capabilities of MLLM4PUE.
-Pathology remains the gold standard for diagnosing a wide range of diseases [42, 32, 8].
-With rapid advancements in artificial intelligence (AI), computational pathology has emerged as a promising field, using AI to address challenges across various tasks such as cancer diagnosis [29, 39], metastasis detection [44, 22], mutation prediction [37, 9, 33], and survival analysis [41, 40, 45].
-These achievements, however, rely on task-specific models trained on large, well-labeled datasets.
-Given the hundreds of tumor types cataloged in the WHO classification111tumourclassification.iarc.who.int/ and the labor-intensive nature of data collection and annotation [12], developing separate, high-performing models for each pathology task is impractical, especially for rare diseases with limited data [24, 30].
-Furthermore, while many efforts focus on image data, pathology often involves natural language, including diagnostic reports and research publications [6, 7, 31, 14].
-Thus, there is a growing need to shift toward task-agnostic models that generate universal multimodal embeddings capable of supporting a broad range of pathology tasks.
-In natural image processing, recent studies such as CLIP [34] have demonstrated remarkable success by aligning visual and language embeddings through contrastive learning, achieving strong performance across diverse downstream tasks.
-Building on these advances, several studies [30, 16, 38, 17, 31] have adapted CLIP-based models for pathology, training on extensive image-text pairs from public sources to produce multimodal embeddings that enable zero-shot learning across various pathology tasks.
+Pathology remains the gold standard for diagnosing a wide range of diseases [41, 32].
+With advancements in artificial intelligence (AI), digital pathology has emerged as a promising field, using AI to address challenges across various tasks such as cancer diagnosis [29, 38], metastasis detection [43, 21], mutation prediction [8, 33], and survival analysis [40, 39, 44].
+These achievements, however, rely on models trained on large, well-labeled datasets.
+Given the hundreds of tumor types cataloged in the WHO classification111tumourclassification.iarc.who.int/ and the labor-intensive nature of data collection and annotation [11], developing separate models for each pathology task is impractical, especially for rare diseases with limited data [23, 30].
+Furthermore, while many efforts focus on image, pathology often involves text, including diagnostic reports and research papers [6, 7, 31, 13].
+Thus, there is a growing need to shift toward task-agnostic models that generate universal multimodal embeddings capable of supporting different pathology tasks.
+Due to the success of CLIP [34] model, several studies [30, 15, 37, 16, 31] have adapted CLIP-based models for pathology, training on extensive image-text pairs to extract embeddings that enable zero-shot evaluation across various pathology tasks.
 However, these methods encounter two significant challenges.
 First, CLIP [34] processes images and text independently, limiting its ability to capture complex interactions between modalities [18].
+Additionally, CLIP [34] can only handle text inputs with a maximum length of 74 tokens, which restricts its ability to leverage longer and more detailed textual descriptions.
 Second, evaluations of these models have mostly centered on isolated tasks, such as classification or retrieval [30, 37, 16], which separate image and text components and do not fully assess multimodal integration capabilities.
-Moreover, inconsistencies in datasets and sample selection criteria across studies lead to non-reproducible and incomparable results, emphasizing the need for a more integrated approach and a standardized benchmarking framework to advance multimodal embeddings in pathology.
-To tackle the first challenge, we propose MLLM4PUE, a framework that uses a Multimodal Large Language Model (MLLM) backbone to generate universal multimodal embeddings for pathology.
-To the best of our knowledge, this is the first use of an MLLM backbone in this field, representing a significant advancement over previous CLIP-based methods.
-Unlike CLIP [34], which processes images and text separately, MLLM4PUE employs a transformer-based architecture to fully integrate these modalities, allowing it to learn from complex image-text relationships.
-This fusion enhances the model’s adaptability and effectiveness across a range of pathology tasks, making it particularly suited for generalizing to diverse pathology scenarios [28, 27].
+Moreover, inconsistencies in datasets and sample selection criteria across studies lead to non-reproducible and incomparable results.
+These challenges emphasize the need for a more integrated model and a standardized benchmarking framework to advance multimodal embeddings in pathology.
+To tackle the first challenge, we propose MLLM4PUE, a framework that uses a MLLM to generate universal multimodal embeddings for pathology.
+Unlike CLIP [34], which processes images and text separately, MLLM employs a transformer-based architecture to fully integrate these modalities, allowing it to learn from complex image-text relationships [28, 27].
+Moreover, MLLM can handle longer textual inputs compared to CLIP, overcoming the 74-token limitation and enabling richer contextual understanding, which is important in pathology area.
+To harness the capability of MLLM in generating embeddings, we take the following steps.
+First, we employ summarization prompts to guide the MLLM to output embeddings.
+Next, we fine-tune MLLM with a LoRA module using pathology image-text pairs, adapting it specifically for the pathology domain.
 To address the second challenge, we introduce the Pathology Multimodal Embedding Benchmark (PMEB) to establish a standardized evaluation platform for multimodal embeddings in pathology.
@@
 Overall, our contributions are summarized as follows:
-  • We introduce MLLM4PUE, the first framework utilizing MLLMs to generate universal multimodal embeddings in pathology.
-This framework integrates multimodal information within a unified model, converting any combination of images and text into robust embeddings.
+  • We introduce MLLM4PUE, a framework converting pathology images and texts into robust embeddings within a unified model.
+This is the first time to explore the possibility of re-purposing MLLM for extracting embeddings in pathology area.
   • We provide PMEB, a comprehensive benchmark for evaluating multimodal embeddings in pathology, encompassing a broad range of tasks to support robust and generalizable assessment of these embeddings.
-  • MLLM4PUE effectively represents multimodal information, consistently surpassing the performance of all baseline models across all evaluated tasks, as demonstrated by the quantitative results in Fig.
+  • MLLM4PUE effectively represents multimodal information, consistently surpassing the performance of baseline models across amounts of evaluated tasks, as demonstrated by the quantitative results in Fig.
 1.
-  • Our approach offers new insights into computational pathology by adopting MLLMs for universal embeddings.
-Future research can build on this work to unify and streamline the development of models for diverse pathology tasks.
-Figure 2: Overview of the MLLM4PUE framework.
-(a) Illustration of pathology image-text contrastive learning with our proposed method.
-With differents prompts, MLLM can generate embeddings for multimodal inputs.
-Image and text embeddings are aligned using a contrastive loss.
-(b) Zero transfer for classification.
-(c) Zero transfer for composed retrieval.
-2 Related work
-2.1 Multimodal embedding learning
-Creating effective multimodal embeddings remains a challenging area of research.
-The CLIP model [34] addresses this by employing separate encoders for images and text, aligning them in a common space through contrastive learning.
-This approach has influenced many subsequent models, such as BLIP [26] and CoCa [43].
-While these models excel in a variety of tasks, the use of distinct encoders restricts their capacity to fully integrate visual and textual data, which is a key requirement for tasks that involve combined visual-language inputs, such as search queries involving both images and text.
-With the rise of large language models (LLMs), Multimodal Large Language Models (MLLMs) have extended LLMs to process multiple data modalities, achieving notable progress in understanding and reasoning across diverse input types [28, 27, 25].
-Although MLLMs exhibit strong capabilities in interpreting multimodal content and following complex instructions, there is still limited research on their application in creating effective multimodal embeddings.
-The recent E5-V [19] model fine-tunes MLLMs with text-based data to improve embedding quality.
-Building on this, we have adapted the approach for pathology by using domain-specific image-text data, resulting in substantial performance improvements.
-2.2 Multimodal embeddings for computational pathology
-In computational pathology, recent studies have adopted the CLIP [34] architecture and training methodology, leveraging paired image-text data within a contrastive learning framework to align similar image-text embeddings while separating dissimilar ones [30, 16, 38, 17, 18].
-For instance, PLIP [16] fine-tunes CLIP on large collections of image-text pairs sourced from platforms like Twitter and other public datasets.
-Similarly, PathCLIP [38] and QuiltNet [17] scale up pathology-specific data for fine-tuning, while CONCH [31] employs over 1.17 million image-caption pairs, making it the largest dataset used for training to date.
-Despite these advancements, these methods fine-tune either CLIP [34] or CoCa [43] using pathology image-caption pairs and rely on separate encoders for processing images and text.
-In contrast, we propose an MLLM-based approach to capture universal multimodal embeddings for pathology by integrating image and text data within a unified model.
-This integrated framework enhances performance across a wide range of computational pathology tasks, offering a more holistic and effective solution.
-3 Methodology
-In this section, we first introduce the Pathology Multimodal Embedding Benchmark (PMEB), a comprehensive benchmark for evaluating multimodal embeddings in diverse pathology tasks.
-Then, we present the MLLM4PUE framework, which integrates multimodal embeddings, contrastive learning, and zero-shot transfer methods to align and adapt pathology image-text data effectively, as illustrated in Fig.
-2.
-3.1 Pathology multimodal embedding enchmark
+  • Our approach offers new insights into digital pathology by adopting MLLM for universal embeddings.
+Future research can build on this work to unify the development of models for diverse pathology tasks.
 Meta-Task Data source Query Target Original Tasks Original samples Selected samples
@@
 Classification Bach [2] I T Breast tissue 399 399
-CRC-100k [20] I T Colorectal cancer 100,000 1,500
+CRC-100k [19] I T Colorectal cancer 100,000 1,000
+Databiox [19] I T Colorectal cancer 922 922
 Digestpath [9] I T Signet ring cell 455 455
 Digestpath [9] I T Colonoscopy tissue 660 660
-Kathercolon [21] I T Colorectal cancer tissue 7,180 1,500
-LC25000 [4] I T Colon adenocarcinoma 10,000 1,500
-LC25000 [4] I T Lung adenocarcinoma 15,000 1,500
+Kathercolon [20] I T Colorectal cancer tissue 7,180 1,000
+LC25000 [4] I T Colon adenocarcinoma 10,000 1,000
+LC25000 [4] I T Lung adenocarcinoma 15,000 1,000
 Osteo [3] I T Osteosarcoma 1,144 1,144
-Renalcell [5] I T Renal tissue 36,687 1,500
-Sicap [36] I T Gleason pattern 12,081 1,500
-Skincancer [23] I T Skin cancer tissue 129,369 1,500
-Wsss4luad [15] I T LUAD tissue 4,693 1,500
+Renalcell [5] I T Renal tissue 36,687 1,000
+Sicap [36] I T Gleason pattern 12,081 1,000
+Skincancer [22] I T Skin cancer tissue 129,369 1,000
+Wsss4luad [14] I T LUAD tissue 10,091 1,000
 Table 1: The statistics of our PMEB benchmark: 15 original tasks from 14 datasets. I and T denote image and text, respectively.
-We propose PMEB (Pathology Multimodal Embedding Benchmark), a comprehensive benchmark designed to evaluate multimodal embeddings in pathology across a diverse range of tasks.
-These tasks include cancer tissue classification [2, 4, 23], Gleason pattern grading [36], and cell identification [10], among other tasks.
-PMEB comprises 15 original tasks from 14 datasets, organized into three meta-tasks: retrieval, classification, and composed retrieval.
+2 Related Work
+Vision-language Models in Digital Pathology.
+In digital pathology, recent studies have adopted the CLIP [34] model, leveraging paired image-text data within a contrastive learning framework to align similar image-text embeddings while separating dissimilar ones [30, 15, 37, 16, 17].
+For instance, PLIP [15] fine-tunes CLIP on large collections of image-text pairs sourced from platforms like Twitter and other public datasets.
+Similarly, PathCLIP [37] and QuiltNet [16] scale up pathology-specific data for training.
+CONCH [31] employs over 1.17 million image-caption pairs to fine-tune CoCa [42] model.
+Despite these advancements, these methods fine-tune either CLIP [34] or CoCa [42] using pathology image-caption pairs and rely on separate encoders for processing images and text.
+In contrast, we propose an MLLM-based framework to capture universal multimodal embeddings for pathology by integrating image and text data within a unified model.
+This integrated framework enhances performance across a wide range of digital pathology tasks, offering a more holistic and effective solution.
+Multimodal Embedding Learning.
+Creating effective multimodal embeddings remains a challenging area of research.
+CLIP  [34] addresses this by employing separate encoders for images and text, aligning them in a common space through contrastive learning.
+This approach has influenced many subsequent models, such as BLIP [25] and CoCa [42].
+While these models excel in a variety of tasks, the use of distinct encoders restricts their capacity to fully integrate visual and textual data, which is a key requirement for tasks that involve combined visual-language inputs, such as search queries involving both images and text.
+With the rise of large language models (LLMs), MLLMs have extended LLMs to process multiple data modalities, achieving notable progress in understanding and reasoning across diverse input types [28, 27, 24].
+Although MLLMs exhibit strong capabilities in interpreting multimodal content and following complex instructions, there is still limited research on their application in creating embeddings.
+The recent work E5-V [18] fine-tunes an LLM with summarization prompts and text-only data to extract embeddings, then integrates a vision module to obtain multimodal embeddings for zero-shot multimodal retrieval tasks.
+However, for pathological adaptation, E5-V [18] requires a domain-specific LLM pretrained on pathology instruction data, which demands detailed and standardized annotations.
+In contrast, our framework, MLLM4PUE can fine-tune MLLM with a LoRA module using existing pathology image-text pairs, bypassing these prohibitive requirements.
+Figure 2: Overview of the MLLM4PUE framework.
+(a) Illustration of pathology image-text contrastive learning with our proposed method.
+With differents prompts, MLLM can generate embeddings for multimodal inputs.
+Image and text embeddings are aligned using a contrastive loss.
+(b) Zero transfer for classification.
+(c) Zero transfer for composed retrieval.
+3 Methodology
+In this section, we first introduce the Pathology Multimodal Embedding Benchmark (PMEB), a comprehensive benchmark for evaluating multimodal embeddings in diverse pathology tasks.
+Then, we present the MLLM4PUE framework, which integrates multimodal embeddings, contrastive learning, and zero-shot transfer methods to align and adapt pathology image-text data effectively, as illustrated in Fig.
+2.
+3.1 Pathology Multimodal Embedding Benchmark
+We propose PMEB, a comprehensive benchmark designed to evaluate pathology multimodal embeddings.
+PMEB include cancer tissue classification [2, 4, 22], Gleason pattern grading [36], and cell identification [9], among other tasks.
+As shown in Table 1, PMEB comprises 15 original tasks from 14 datasets, organized into three meta-tasks: retrieval, classification, and composed retrieval.
 All tasks are reformulated to assess multimodal embeddings, where the model receives an instructional prompt and a query, which may consist of text, images, or both, and selects the correct target from a set of options.
@@
 This composed retrieval task, specifically designed for multimodal embedding evaluation, incorporates integrated visual and language inputs and has not been applied in previous methods.
-Detailed dataset statistics are provided in Table  1.
-Due to the large size of certain classification datasets, we have standardized their sizes.
-For datasets containing more than 1,500 samples, we employ the CONCH model [31] to calculate similarity scores between each image and its corresponding class label text.
-We then select the top 500 images with the highest scores, the bottom 500 with the lowest scores, and a random sample of 500 from the remaining data, ensuring the original category distribution is preserved.
+Due to the large size of some datasets, we standardize their sizes.
+For datasets containing more than 1,000 samples (except Osteo [3]), we choose a random sample of 1,000, ensuring the original category distribution is preserved.
 For the composed retrieval task, we select only open-ended questions from the data source, where answers are more complex than a simple “yes” or “no”.
 Additionally, since there is an overlap between questions and answers in the VQA data source, we utilize ChatGPT [1] for data cleaning to ensure clarity and accuracy in task evaluation.
-Further details and examples from each dataset within PMEB are provided in Supplementary Materials.
-Figure 3: Illustration of the MLLM4PUE framework for generating multimodal embeddings using an MLLM.
-The image is processed through an Image Encoder and projection layer, with a prompt guiding the MLLM to distill the image content into a single-word embedding.
-The final embedding $h$ is extracted from the last token as the unified multimodal representation.
+Further details of each dataset within PMEB are provided in Supplementary Materials.
 3.2 Multimodal Embeddings with MLLM
@@
 Unlike CLIP, which directly generates embeddings for multimodal inputs, MLLMs are inherently generative models and require a different approach to produce embeddings.
-Inspired by E5-V [19], our solution uses a prompt-based strategy specifically designed to guide MLLMs in embedding multimodal data.
+We uses a prompt-based strategy specifically designed to guide MLLMs in embedding multimodal data.
 […]
@@
 where $\tau$ is a temperature parameter and $cos(h_{vi},h_{ti})$ and $cos(h_{ti},h_{vi})$ represent the cosine similarity in both directions in contrastive learning.
+Figure 3: Illustration of the MLLM4PUE framework for generating multimodal embeddings using an MLLM.
+The image is processed through an Image Encoder and projection layer, with a prompt guiding the MLLM to distill the image content into a single-word embedding.
+The final embedding $h$ is extracted from the last token as the unified multimodal representation.
 3.4 Zero-shot Transfer Evaluation
@@
 The similarity between these embeddings is calculated as outlined in Section 3.3, and test image labels are assigned based on the highest similarity scores.
-To account for variability in templates, we randomly select one from the 21 templates in CONCH [31] and repeat this process 100 times for each dataset.
-The templates are exhibited in Supplementary Materials.
 […]
@@
 Bold values indicate the best performance.
+Same color represents models trained on the same dataset.
 4 Experiments
-Figure 4: Comparison of zero-shot patch classification performance across different models, measured by weighted F1 metric.
-Each box represents 100 data points, with each point indicating the performance on one template.
-The upper, middle, and lower lines of each box denote the first, median, and third quartiles, respectively.
-MLLM4PUE-O, MLLM4PUE-P, and MLLM4PUE-Q denote our model pre-trained on Openpath [16], PathCap [38], Quilt1M [17], respectively.
-4.1 Training Datasets
+Training datasets.
 We leverage three public datasets, Openpath [15], PathCap [37] and Quilt1M [16], to construct a robust training foundation for pathology image-text pairs.
@@
 Through these careful preparations, we compile an extensive training dataset consisting of 593,838 image-text pairs.
-4.2 Baselines
-We evaluate the performance of our proposed MLLM4PUE, against several recent SOTA approaches across various downstream tasks.
-Specifically, we compare with three currently developed CLIP-based methods: PLIP [16], PathCLIP [38], and QuiltNet [17].
-We also compare our method with E5-V [19], a contemporary model that leverages vision-language models for multimodal embedding tasks.
-It proposes a single-modality training approach, where the model is trained exclusively on text pairs.
-4.3 Implementation Details and Metrics
-We use LLaVA-NeXT [25] as our backbone MLLM.
-To save the GPU memory, we use QLoRA [11] and gradient checkpointing with DeepSpeed ZeRO-2, and all the images are resized to 336 × 336 pixels.
-The temperature is set to 0.02.
-All experiments are run on six H100 GPUs.
+Baseline Methods and Metrics.
+We evaluate the performance of our proposed MLLM4PUE against several recent baseline models across various downstream tasks.
+Specifically, we compare it with three pathology CLIP-based models: PLIP [15], PathCLIP [37], and QuiltNet [16], all of which are trained on the same dataset as our method.
+Additionally, we compare our approach with E5-V [18], a method that fine-tunes an LLM with summarization prompts and text-only data to extract embeddings, then integrates a vision module to obtain multimodal embeddings for zero-shot multimodal retrieval tasks.
+For zero-shot classification tasks, we also compare our method with CONCH [31], a model trained on over 1.17 million private image-caption pairs, whereas our method is trained on fewer than 0.6 million pairs.
 To evaluate the performance of retrieval and composed retrieval tasks, we use the Recall@K metric, which quantifies the proportion of relevant items contained within the top K retrieved results.
@@
 The weighted F1 is calculated by averaging the F1 scores for each class, with each class’s score weighted by its frequency of occurrence.
+Implementation Details.
+Our framework is implemented in Pytorch and we use LLaVA-NeXT-8B [24] as our backbone MLLM.
+To save the GPU memory, we use QLoRA [10] and gradient checkpointing with DeepSpeed ZeRO-2 to fine-tune the MLLM, and all the images are resized to 336 × 336 pixels.
+The temperature is set to 0.02.
+All experiments are run on six H100 GPUs with a gradient accumulation batch size of 144 and a learning rate of 4e-4.
+Model Bach
+Data
+biox
+Digest
+Digest
+CRC
+100k
+Kather
+7k
+LC
+colon
+LC
+lung
+Osteo
+Renal
+cell
+Sicap
+Skin
+cancer
+Wsss
+4luad
+E5-V [18] 0.100 0.162 0.197 0.049 0.224 0.206 0.333 0.209 0.300 0.051 0.230 0.017 0.132
+CONCH [31] 0.511 0.413 0.835 0.824 0.680 0.625 0.890 0.650 0.382 0.393 0.359 0.593 0.772
+PLIP [15] 0.124 0.194 0.696 0.252 0.503 0.504 0.697 0.743 0.554 0.413 0.080 0.311 0.394
+MLLM4PUE-O 0.561 0.437 0.784 0.480 0.642 0.509 0.958 0.744 0.581 0.507 0.284 0.411 0.636
+PathCLIP [37] 0.528 0.282 0.785 0.707 0.611 0.492 0.937 0.887 0.668 0.359 0.267 0.332 0.648
+MLLM4PUE-P 0.577 0.388 0.799 0.900 0.763 0.783 0.937 0.886 0.751 0.543 0.502 0.504 0.754
+QuiltNet [16] 0.403 0.318 0.491 0.564 0.564 0.591 0.884 0.801 0.578 0.435 0.322 0.467 0.364
+MLLM4PUE-Q 0.437 0.434 0.868 0.851 0.686 0.668 0.886 0.866 0.613 0.516 0.440 0.507 0.758
+E5-V [18] 0.251 0.324 0.367 0.169 0.299 0.306 0.500 0.348 0.385 0.166 0.337 0.053 0.267
+CONCH [31] 0.559 0.444 0.832 0.802 0.703 0.644 0.890 0.665 0.439 0.447 0.405 0.607 0.783
+PLIP [15] 0.251 0.306 0.708 0.253 0.532 0.541 0.721 0.749 0.551 0.488 0.179 0.290 0.450
+MLLM4PUE-O 0.569 0.448 0.793 0.429 0.666 0.546 0.958 0.775 0.624 0.518 0.426 0.453 0.671
+PathCLIP [37] 0.556 0.386 0.784 0.666 0.610 0.489 0.937 0.888 0.698 0.350 0.362 0.378 0.688
+MLLM4PUE-P 0.599 0.386 0.812 0.892 0.767 0.784 0.937 0.886 0.744 0.576 0.509 0.496 0.760
+QuiltNet [16] 0.441 0.413 0.628 0.503 0.607 0.586 0.904 0.802 0.588 0.434 0.338 0.458 0.449
+MLLM4PUE-Q 0.479 0.441 0.866 0.849 0.695 0.661 0.887 0.872 0.609 0.494 0.444 0.520 0.761
+Table 3: Performance comparison with baseline methods on zero-shot classification tasks, reporting weighted F1 (wF1) (top) and accuracy (bottom) metrics.
+MLLM4PUE-O, MLLM4PUE-P, and MLLM4PUE-Q denote our model pre-trained on Openpath [15], PathCap [37], Quilt1M [16], respectively.
+Bold values indicate the best performance.
+Same color represents models trained on the same dataset.
 5 Results
+5.1 Zero-shot Cross-modal Retrieval
+We first evaluate our methods on the retrieval task on Arch-PubMed [12] and Arch-book datasets [12], where the query is an image or text, with the target being the corresponding paired text or image.
+Recall@k metric is employed to assess the model’s ability to retrieve relevant pathology image-text pairs among top k results.
+As shown in Table 2, our models consistently outperform the baseline models, especially at higher recall values.
+Notably, in the image-to-text (i2t) task on Arch-book dataset [12], our MLLM4PUE trained on Openpath dataset [15] achieves a Recall@5 of 0.185, indicating an approximate 9% improvement over PLIP [15].
+Moreover, our PathCap-pretrained model surpasses the performance of PathCLIP [37] on both benchmark datasets.
+A similar enhancement is observed in the text-to-image (t2i) retrieval task.
+For instance, on the Arch-PubMed dataset [12], our model achieves a Recall@10 of 0.193, marking a 10% increase in performance.
+These results collectively demonstrate that our framework excels at capturing multimodal embeddings more effectively than CLIP-based models with the same training data.
+Furthermore, it is important to note that the E5-V model [18] yields significantly lower recall scores across all tasks and datasets due to the lack of fine-tuning with pathology-specific data.
 Model Quilt-VQA [35] Quilt-VQA-RED [35]
@@
 MLLM4PUE-Q 0.475 0.598 0.823 0.675 0.754 0.948
-Table 3: Zero-shot composed retrieval performance on the Quilt-VQA and Quilt-VQA-RED datasets, with Recall metrics reported at various thresholds.
+Table 4: Zero-shot composed retrieval performance on the Quilt-VQA [35] and Quilt-VQA-RED [35] datasets, with Recall metrics reported at various thresholds.
 MLLM4PUE-O, MLLM4PUE-P, and MLLM4PUE-Q denote our model pre-trained on Openpath [15], PathCap [37], Quilt1M [16], respectively.
-Bold values indecate the best performance.
-Prompt-tuning Prompt Arch-PubMed [13] Arch-book [13]
-i2t t2i i2t t2i
-R@5 R@10 R@5 R@10 R@5 R@10 R@5 R@10
-no no prompt 0.086 0.135 0.087 0.137 0.147 0.213 0.149 0.226
-no “in one word” 0.092 0.137 0.095 0.142 0.152 0.236 0.158 0.233
-yes no prompt 0.091 0.139 0.091 0.141 0.152 0.233 0.161 0.246
-no “in one word” 0.094 0.143 0.102 0.149 0.157 0.243 0.165 0.258
-MLLM4PUE 0.098 0.155 0.109 0.162 0.169 0.253 0.177 0.274
-Table 4: Impact of prompt-tuning on two retrieval datasets.
-All models are trained on the Openpath dataset [16] for 1000 steps.
-Task i2t and t2i denote image-to-text and text-to-image retrieval.
-R@5 and R@10 represent Recall@5 and Recall@10, respectively.
 Bold values indicate the best performance.
+Same color represents models trained on the same dataset.
+5.2 Zero-shot Patch Classification
+We then evaluate the performance of zero-shot classification on patch-level pathology images across 13 classification tasks from our PMEB.
+In this downstream task, each query is an image, and the target is an extended sentence of the class name as described in Section 3.4.
+Table 3 illustrates that our MLLM4PUE consistently surpasses baseline models on most datasets.
+Notably, with the same training data, MLLM4PUE-O outperforms PLIP [15] across all datasets.
+Similarly, MLLM4PUE-Q achieves the best performance compared to QuiltNet [16], demonstrating the effectiveness of our pretraining strategy on the same training dataset.
+While CONCH [31] attains the highest score on SkinCancer [22] and Wsss4luad [14], its superior performance stems from utilizing a much larger private dataset (1.16M).
+In contrast, our model, MLLM4PUE-P, outperforms all baseline models across seven datasets with only 0.2 million training samples, and is only 0.02 less than CONCH [31] on the Wsss4luad dataset [14].
+This underscores the advantages of leveraging vision-language alignment in pathology representation learning.
+These findings collectively highlight the effectiveness of our approach, especially considering that our models were trained on fewer samples, further demonstrating their data efficiency in pathology-specific zero-shot classification.
+The accuracy metric shows a similar trend, where MLLM4PUE-P outperforms the baselines on most datasets.
+5.3 Zero-shot Composed Retrieval
+[…]
+5.4 Visualization of Modality Gap
+To illustrate the modality gap, we employed t-SNE visualization to compare the embeddings of our MLLM4PUE-P model with those of PathCLIP [37], both trained on the Patchcap dataset [37].
+As shown in Fig 4, the qualitative analysis clearly shows that the embeddings produced by MLLM4PUE-P form tighter clusters in the low-dimensional space, whereas those from PathCLIP [37] are more dispersed.
+This difference reflects a significant improvement in cross-modal alignment of our method.
+Figure 4: Distribution of image and text embeddings on Arch-PubMed [35] and Arch-Book [35] dataset.
+To further quantify the modality gap, we adopted the metric $||\Delta||_{gap}$ as proposed in Liang et al.
+[26].
+Our experiments reveal that, on the Arch-PubMed dataset [35], our method reduces the gap from 0.9284 to 0.2587, and on the Arch-Book dataset [35], from 1.0408 to 0.3739.
+These results provide strong evidence that our model can significantly narrow the gap between image and text embeddings.
+5.5 Ablation Studies
 Model Quilt-VQA [12] Quilt-VQA-RED [12]
@@
 Bold values indicate the best performance.
-5.1 Zero-shot Cross-modal Retrieval
-We first evaluate our methods on the retrieval task, where the query is an image or text, with the target being the corresponding paired text or image.
-Recall@k metric is employed to assess the model’s ability to retrieve relevant pathology image-text pairs among top k results.
-Our MLLM4PUE models are pretrained on Openpath [16], PatchCap [38], and Quilt1M [17].
-As shown in Table 2, our models consistently outperform the baseline models, especially at higher recall values.
-Notably, in the image-to-text (i2t) task on Arch-book dataset [13], our MLLM4PUE trained on Openpath dataset [16] achieves a Recall@5 of 0.185, indicating an approximate 9% improvement over PLIP [16].
-Moreover, our PathCap-pretrained model surpasses the performance of PathCLIP [38] on both benchmark datasets.
-A similar enhancement is observed in the text-to-image (t2i) retrieval task.
-For instance, on the Arch-PubMed dataset [13], our model achieves a Recall@10 of 0.193, marking a 10% increase in performance.
-These results collectively demonstrate that our framework excels at capturing multimodal embeddings more effectively than CLIP-based models with the same training data.
-Furthermore, it is important to note that the E5-V model [19], although sharing the same underlying architecture as ours, yields significantly lower recall scores across all tasks and datasets due to the lack of fine-tuning with pathology-specific data.
-5.2 Zero-shot Patch Classification
-We then evaluate the performance of zero-shot classification on patch-level pathology images.
-Specifically, each query is an image, and the target is an extended sentence of the class name, constructed using a template outlined in sec 3.4.
-Following CONCH [31], we conduct 100 experimental trials for each dataset, randomly selecting one template from a set of 21 templates for each trial.
-The detailed descriptions are available in the Supplementary Material.
-We present the distribution of weighted F1 scores (wF1) from these 100 experiments in Fig.
-4, where the top, middle, and bottom lines of each box depict the first quartile, median, and third quartile of the wF1 score distribution, respectively.
-Each point in the figure indicates a single template’s performance.
-Fig.
-4 illustrates that our MLLM4PUE consistently surpasses baseline models trained on identical datasets.
-For instance, in the “DigestpathSignet” task with the Openpath dataset [16], which focuses on identifying signet ring cells, MLLM4PUE-O achieves a median wF1 score of 0.665, marking an improvement of over 30% compared to PLIP’s median of 0.329.
-When the training data changes to PathCLIP [38], MLLM4PUE-P also shows considerable advancements, achieving a median score of 0.854 compared to PathCLIP’s 0.361.
-This implies a performance increase of roughly 50%.
-Similarly, using the Quiltnet dataset [17], MLLM4PUE-Q excels in the “Sicap” dataset, reaching a median wF1 of 0.402, far exceeding Quiltnet’s 0.164.
-In the “Osteo” dataset, MLLM4PUE-Q achieves a median wF1 score of 0.671, significantly outclassing Quiltnet with an improvement exceeding 40%.
-Moreover, when compared with E5V, MLLM4PUE demonstrates superior performance, delivering higher wF1 scores even with the same architectural foundation.
-These results confirm the superiority of MLLM4PUE in enhancing zero-shot classification performance through effective multimodal data integration, establishing it as a robust choice for various pathology image classification tasks.
-5.3 Zero-shot Composed Retrieval
-We further evaluate the performance of composed retrieval on two datasets, where the task involves using an image and a question as the query to retrieve the correct answer.
-The Recall@k metric is employed to assess how well the model can identify the correct answer within the top k candidates.
-As shown in Table 3, our proposed MLLM4PUE demonstrates significant improvements compared to baseline methods, even when trained with the same pre-training data.
-Specifically, on the Quilt-VQA dataset [35], MLLM4PUE achieves an increase in Recall@5 by approximately 4%, 3%, and 30% over PLIP [16], PathCLIP [38], and QuiltNet [17], respectively.
-Similarly, on the Quilt-RED-VQA dataset [35], MLLM4PUE pre-trained on Quilt1M [17], has achieved a Recall@5 of 0.675, which represents the best performance among all baseline methods.
-Interestingly, the E5-V model [19] exhibits superior performance compared to all CLIP-based models, despite lacking fine-tuning on pathology-specific data.
-This demonstrates the robust capacity of MLLM backbone in multimodal embedding fusion, highlighting its proficiency in handling diverse data inputs effectively.
-5.4 Ablation studies
-5.4.1 Impact of Prompts
-To assess the impact of different prompt strategies, we conduct a comparative analysis between our MLLM4PUE and two alternative approaches: no prompt - where images or text are directly input into the model without prompts, and no “in one word” - employing a similar prompt to our method but omitting the phrase “in one word”.
-Table 4 presents the performance of these methods, both with and without fine-tuning.
-All models are trained on Openpath dataset [16] for 1000 steps.
-For models employing fine-tuning, each method is adjusted with its respective prompts.
-In contrast, the no fine-tuning scenario involves simply altering the prompt for the already finely-tuned MLLM4PUE model.
-Our method shows significant improvements on zero-shot retrieval tasks compared to the no prompt and no “in one word”.
-For the Arch-PubMed dataset, MLLM4PUE attains Recall@5 scores of 0.098 and 0.109 for image-to-text and text-to-image tasks, respectively, marking the best performance in those categories.
-Similarly, in the Arch-book dataset, MLLM4PUE achieves the top Recall@5 and Recall@10 scores, indicating superior retrieval effectiveness.
-These results highlight the significant enhancements provided by our prompt, emphasizing its role in facilitating more accurate zero-shot retrieval in complex datasets by better aligning multimodal inputs.
-5.4.2 Fusion Method in Composed Retrieval
-This experiment is designed to highlight the effectiveness of our MLLM4PUE in modality fusion while exploring the limitations of CLIP in handling composed retrieval tasks.
-As shown in Table 5, the results reveal that QuiltNet [17], a CLIP-based model, shows little difference in performance between using only the question text and integrating image embeddings.
+Robustness of PMEB.
+Since some datasets in our PMEB are super large, we choose a random sample of 1,000.
+To assess the robustness of our benchmark, we conduct an ablation study by computing the wf1 scores using 1,000, 3,000, 5,000, and all available samples for each dataset.
+As shown in Figure 5, the performance change between 1,000 and the full dataset remains within a small margin across all datasets, confirming that using only 1,000 samples provides a reliable approximation of the complete dataset’s performance.
+Given that the performance difference is typically within 0.01 to 0.04 across different sample sizes, using 1,000 samples significantly reduces computational costs while maintaining reliable performance.
+This makes PMEB highly efficient for large-scale experiments without sacrificing meaningful evaluation.
+We also report the result of accuracy metric and the results in CONCH [31] and PathCLIP [37] model in the Supplementary Materials, which shows the same trend.
+Figure 5: Ablation study on the sample numbers of our benchmark on our MLLM4PUE-P model. We computed the wF1 scores using 1,000, 3,000, 5,000, and all available samples for each dataset.
+Embedding Fusion Methods.
+This experiment is designed to validate the quality of our benchmark and highlight the effectiveness of our MLLM4PUE in modality fusion compared to CLIP-based model in composed retrieval tasks.
+As shown in Table 5, MLLM4PUE outperforms QuiltNet [16] in every setting.
+First, comparing “only question” results, our MLLM4PUE shows significant improvement compared with QuiltNet [16], while both models demonstrate substantial performance gains when incorporating visual information via “add embedding”.
+This consistent pattern across both models confirms that our benchmark has been properly curated, requiring true integration of visual and textual information rather than allowing models to infer answers solely from questions.
+Second, the results reveal that QuiltNet [16], a CLIP-based model, shows little difference between using only the question text and integrating image and text embeddings.
 This indicates that CLIP’s method lacks robust modality integration, as the inclusion of visual information fails to significantly enhance its effectiveness.
-Conversely, our model, MLLM4PUE, demonstrates substantial improvements by effectively fusing image and text information.
-MLLM4PUE outperforms QuiltNet in every setting.
+Conversely, our model MLLM4PUE demonstrates substantial improvements by effectively fusing image and text information.
 The true strength of our approach becomes evident when employing a prompt-based method, leading to significantly higher recall scores.
@@
 6 Conclusion
-In conclusion, our work represents a significant advancement in different downstream tasks through MLLM4PUE, a novel framework leveraging MLLMs to achieve universal multimodal embeddings.
-By effectively integrating image and text data, our model addresses the limitations of previous CLIP-based methods.
+In this paper, we have explored a novel paradigm for applying MLMMs to generate multimodal universal embeddings for a broad range of pathology downstream tasks.
+By applying summarization prompts and inserting lightweight LoRA modules, our proposed framework, MLLM4PUE, can effectively integrates image and text data, and addresses the limitations of previous CLIP-based methods.
 We also introduce PMEB, which provides a standardized and comprehensive benchmark for evaluating multimodal capabilities, promoting reproducibility and comparability in the field.
-This study also paves the way for future research to utilize MLLM-based models to support a wide range of downstream tasks and unify the research direction for foundation models in pathology.
+Comprehensive experiment results across 15 diverse zero-shot retrieval and classification tasks underscore the versatility and effectiveness of our method.
+We anticipate that our framework will provide valuable insights into the pathology vision-language foundation models and inspire further research in this area.